OpenStudy (96mertcank96):

!!!!!!!HELP!!!!!!! Michael's father bought him a 16 foot board to cut into shelves for his bedroom. Michael plans to cut the board into 11 equal size lengths for his shelves. The saw blade that michael will use to cut the board will change the length of the board by -0.125 inches for each cut. How will this affect the total length of the board??

OpenStudy (anonymous):

hint: you must convert inches to feet

OpenStudy (96mertcank96):

Thanks man but can you give the answer and work please?

OpenStudy (anonymous):

actually, you can convert 16 feet to inches

OpenStudy (anonymous):

alright, i subtracted 192 inches (total inch of board) by 11 (pieces) x .125

OpenStudy (anonymous):

then i divided by 11

OpenStudy (anonymous):

and i got about 17.33 inches per board and to answer the question, it will affect the total length of the board by -1.375 inches.

OpenStudy (anonymous):

YOUR WELCOME

OpenStudy (96mertcank96):

so you did 192 - 11= 181 and then you multiplied it by .125 to get 22.625 then you divided by 11?

OpenStudy (anonymous):

11x ( -.125) = - 1.375 inches then 192 - 1.375 = 190.625 then divide 190.625 / 11 = 17.33 inch

OpenStudy (anonymous):

but all it is asking for is the first and second equation.

OpenStudy (anonymous):

you understand?

OpenStudy (96mertcank96):

Ohh thanks man!

OpenStudy (anonymous):

no problem man. give me a trophy brotha

OpenStudy (96mertcank96):

Do you mind helping me with another question? Same problem

OpenStudy (anonymous):

oh i think you already did lol. i'll walk you through your next one. i won't do your homework for you though

OpenStudy (anonymous):

;)

OpenStudy (96mertcank96):

After making his cuts, what will the exact length of each shelf be?

OpenStudy (anonymous):

i told you already :) you must understand what i showed you

OpenStudy (96mertcank96):

but so the exact length of the board is 17.33 inches?

OpenStudy (anonymous):

to be exact its 17.32954545454545454545455454545455454545454554545455454

OpenStudy (anonymous):

inches

OpenStudy (96mertcank96):

Ok but in the first question it says how will this affect the total length of the board?

OpenStudy (anonymous):

i told you already. I gave you all the answers. it seems you don't understand the questions though. I can't help you out or you will never succeed. just try the calculations i did and label your calculations in feet and inches. then you will understand. Good luck buddy.